OpenStudy (anonymous):

Determine whether Rolle's Theorem can be applied to the function on the given interval; if so, find the value(s) of c guaranteed by the theorem. (Enter your answers as a comma-separated list. If Rolle's Theorem does not apply, enter DNE) f(x) = ln(5 − sin(x)) on [0, 2π]

OpenStudy (anonymous):

Do you know the conditions of Rolle's Theorem?

OpenStudy (anonymous):

no i don't !

OpenStudy (anonymous):

The function has to be continuous and differentiable and f(a)=f(b)

OpenStudy (anonymous):

I guess I can join in?

OpenStudy (anonymous):

yeah i do but i tried and didn't get the right answer

OpenStudy (anonymous):

just answered this question earlier on line class?

OpenStudy (anonymous):

Okay so you know rolle's theorem states that if you have a continuous and differentiable function on [a,b] and (a,b) then there exists a number c where f'(c) is equal to zero on that interval.

OpenStudy (anonymous):

So you would differentiate and set y= 0, then solve for x.

OpenStudy (anonymous):

sinc \(sin(0)=\sin(2\pi)=0\) you get \(f(0)=f(2\pi)=\ln(5)\) take the derivative, set it equal to zero and solve for \(x\)

OpenStudy (anonymous):

okay ! thanks for the help and yes online class :(